A1g of an element gives A2g of its oxide.the equivalent mass of the element is?

Grade:12

1 Answers

Arun
5 years ago
Dear student!The exact weight could not be calculated as data is not given, however it may be solved as:Weight of metal = A1 gm ;Weight of oxide = A2 gmOxygen consumed = A2 - A1 gmSince, (A2 - A1) gm of oxygen combines with A1 gm of metal So, 1 gm of oxygen will combine with = A1/ (A2 - A1) gm of metalHence, 8 gm of oxygen will combine with = 8*A1/ (A2 -A1) gm of metalThis 8 *A1/(A2-A1) quantity is the equivalent weight of metal